OpenStudy (anonymous):

Another Proving of Identities: sin^3x+cos^3x/(sinx+cosx) = 1-sinxcosx ? How ?

OpenStudy (loser66):

a^3 + b^3 =?

OpenStudy (anonymous):

um? sir? sorry can't follow.

zepdrix (zepdrix):

lol, you and your `sir`'s :p

OpenStudy (anonymous):

???

OpenStudy (loser66):

zepdrix will answer. hehehe.

OpenStudy (anonymous):

ok.

zepdrix (zepdrix):

There are formulas for the sum and difference of `cubes`. They don't come up very often but it's still worth remembering them I think. :)\[\large a^3+b^3=(a+b)(a^2-ab+b^2)\]

OpenStudy (anonymous):

oh yes yes I remember that.

zepdrix (zepdrix):

\[\large \sin^3x+\cos^3x=?\]

OpenStudy (anonymous):

1

zepdrix (zepdrix):

No no no silly :3 that's \(\large \sin^2x+\cos^2x\)

OpenStudy (anonymous):

oh sorry. (_ _)

OpenStudy (anonymous):

(sinx+cosx)(sin^2x+2sinxcosx+cos^2x) ?

OpenStudy (anonymous):

wait the degrees are wrong I think

zepdrix (zepdrix):

Follow the formula buster! :O

zepdrix (zepdrix):

The powers looks correct, its a couple other minor mistakes i see though :U

OpenStudy (anonymous):

(sinx+cosx)(sin^2x-2sinxcosx+cos^2x) ? it's minus

zepdrix (zepdrix):

Yah it's minus :) Also, that 2 shouldn't be there in the middle, on the -2sinxcosx.

OpenStudy (anonymous):

so it's sinxcosx only?

OpenStudy (anonymous):

yup

zepdrix (zepdrix):

Ya, (sinx+cosx)(sin^2x-sinxcosx+cos^2x)

OpenStudy (anonymous):

oh right.:) then how can I prove it?

zepdrix (zepdrix):

So using that identity, it allows us to rewrite our numerator like this, correct? \[\large \frac{\sin^3x+\cos^3x}{(\sin x+\cos x)} \qquad=\qquad \frac{\color{orangered}{(\sin x+\cos x)}(\sin^2x-\sin x \cos x+\cos^2x)}{\color{orangered}{(\sin x+ \cos x)}}\] See any nice cancellations from here? :3 Maybe I made it way too obvious lol

OpenStudy (anonymous):

I get it now

zepdrix (zepdrix):

oh ok c: too quick for me hehe

OpenStudy (anonymous):

thanks sir. :)

OpenStudy (anonymous):

gr* stuff man @zepdrix

OpenStudy (anonymous):

i mean great :)